### v5

# M3GPT: An Advanced Multimodal, Multitask Framework for Motion Comprehension and Generation

Abstract
This paper presents M3GPT, an advanced Multimodal, Multitask framework for Motion comprehension and generation. M3GPT operates on three fundamental principles. The first focuses on creating a unified representation space for various motion-relevant modalities. We employ discrete vector quantization for multimodal conditional signals, such as text, music and motion/dance, enabling seamless integration into a large language model (LLM) with a single vocabulary. The second involves modeling motion generation directly in the raw motion space. This strategy circumvents the information loss associated with a discrete tokenizer, resulting in more detailed and comprehensive motion generation. Third, M3GPT learns to model the connections and synergies among various motion-relevant tasks. Text, the most familiar and well-understood modality for LLMs, is utilized as a bridge to establish connections between different motion tasks, facilitating mutual reinforcement. To our knowledge, M3GPT is the first model capable of comprehending and generating motions based on multiple signals. Extensive experiments highlight M3GPT’s superior performance across various motion-relevant tasks and its powerful zero-shot generalization capabilities for extremely challenging tasks. Project page: https://github.com/luomingshuang/M3GPT.
Figure 1: M3GPT can handle core motion comprehension and generation tasks, including text-to-motion, motion-to-text, music-to-dance, dance-to-music, motion prediction, and motion in-between. The motion sequences within the dashed-line areas are masked in the input.
1 Introduction
Motion comprehension and generation in multimodality are crucial for diverse applications, including AR/VR creation, video games, and virtual reality. Numerous studies [15, 51, 64, 52] focus on motion comprehension, including captioning 3D human motions and generating music from 3D human dances [^1] In this paper, the term ”motion” generally includes ”dance.” We distinguish them when referring to specific tasks or scenes, such as text-to-motion, and music-to-dance.. Recent advancements in AI [14, 44, 42, 48] have paved the way for motion generation, allowing for various control signals including textual descriptions, music pieces, and human poses. A significant shortcoming of most existing works is their focus on single-modality control signals, overlooking the potential for multimodal information integration. More importantly, the comprehension and generation of motions are predominantly studied in isolation. In reality, human motion cognition and communication indispensably require seamless transitions between any motion-relevant modalities. Therefore, it is vital to develop a unified framework for motion comprehension and generation that can efficiently utilize multiple signals simultaneously.
Recent works [12, 62, 63, 58] have shown success in developing a unified multitask motion framework which integrates text-driven and audio-driven motion generation through a single architecture. Employing a large language model (LLM), [60] adeptly handles multimodal control signals, such as text and single-frame pose, to generate consecutive motions. Despite their promising performance in motion generation, these approaches often fall short in comprehending motion. MotionGPT [21], a recent innovation, constructs a unified motion-language model to generate plausible human motions and natural language descriptions through prompt instructions. However, MotionGPT focuses solely on a single non-motion modality, i.e., text. While aligning motion with one additional modality is relatively straightforward, integrating three or more modalities within a single framework and achieving bidirectional alignment among them to cover a broad range of modalities for motion comprehension and generation presents a formidable challenge.
Two main challenges need to be solved for building a unified multimodal framework for motion comprehension and generation. The first is how to create a unified representation space across different motion-relevant modalities. MotionGPT [21] and SpeechGPT [54] separately treat motion and speech as specific language for seamlessly integrating with text. Inspired by these efforts [21, 54], we view both motion and music as distinct forms of language, facilitating better associations with text via LLMs. Specifically, akin to language, we compress raw motion and music into a sequence of discrete semantic tokens. By encoding motion, music, and language within a single vocabulary, we can build a unified representation space across these different modalities. The second is how to model the connections and synergies among various motion tasks. Different motion-relevant tasks are interconnected and can mutually enhance each other. Since text is the most familiar and well-understood modality for LLMs, we propose employing text as a bridge to establish connections between different motion tasks. Specifically, to better learn the complex music-to-dance task where both input and output modalities are unfamiliar to LLMs, we introduce two auxiliary tasks: music-to-text and text-to-dance, aimed at aligning music and dance modalities with the structured text embedding space. This strategy enables us to establish connections and synergies between music-to-dance and text-to-motion tasks, facilitating the alignment and collaboration of text, music, and motion/dance modalities across different tasks.
In this work, we propose a uniform Multimodal, Multitask framework for Motion comprehension and generation, namely M3GPT, that leverages the strong language generation capability of LLMs for unifying various motion-relevant tasks, as depicted in Fig. 1. M3GPT comprises three tires. Firstly, M3GPT is equipped with multimodal tokenizers capable of compressing raw multimodal data, including motion, music, and text, into a sequence of discrete semantic tokens. These discrete representations allow the core LLM to unify motion comprehension and generation in an autoregressive manner, operating at the discrete semantic representation space. Secondly, different from [21, 60] that solely optimize LLM in discrete semantic space, we jointly train LLM and motion de-tokenizer, optimizing LLM in both discrete semantic space and raw continuous motion space. This operation enables the motion-space error signals from de-tokenizer to backpropagate to LLM, enhancing LLM’s ability to generate the details of motion. Thirdly, we construct paired text descriptions for music, and design two auxiliary music-to-text and text-to-dance tasks, which aid in aligning music and dance with the text embedding space. Also, we build up a shared tokenizer for motion and dance data to project them into a shared semantic space. These auxiliary tasks and shared tokenizer establish connections between music-to-dance and text-to-motion, enabling mutual reinforcement.
We employ a multimodal pre-training + instruction-tuning pipeline to train M3GPT, enhancing inter-modal alignment and effectively aligning them with human intent. To our knowledge, M3GPT is the first approach to integrate six core tasks of motion comprehension and generation—text-to-motion, motion-to-text, music-to-dance, dance-to-music, motion prediction, and motion in-between—into a uniform framework. Extensive experiments demonstrate that M3GPT achieves competitive performance across multiple motion-relevant tasks. Additionally, through qualitative results, we demonstrate that M3GPT possesses powerful zero-shot generalization capabilities, e.g., long-term dance generation and music-text conditioned dance generation.
Methods T2M M2T A2D D2A M2M Random M Random T Random A
TM2D[12]
UDE[62]
MotionGPT[60]
MotionGPT[21]
M3GPT (Ours)
Table 1: Comparison of recent multimodal, multitask methods across various motion comprehension and generation tasks. T2M: text-to-motion; M2T: motion-to-text; A2D: music-to-dance; D2A: dance-to-music; M2M: motion-to-motion that includes motion prediction and motion in-between. Random M, Random T, and Random A represent the unconstrained generation of motion, text, and music [^3] Note: in this paper, the term "audio" specifically refers to "music". This designation is adopted to avoid confusion between the initial letter "M" shared by both "music" and "motion," which could lead to ambiguity when these modalities are represented by their initials., respectively.
2 Related Work
Motion comprehension and Generation.  Many existing works focus on studying human appearance, pose, detection, attribute, part parsing and so on [61, 19, 45, 40, 23, 17]. This work focuses on studying human motion, including motion comprehension and motion generation. Motion comprehension involves two core tasks: motion-to-text and dance-to-music. Motion-to-text aims to describe human motion with natural language [37]. For example, recurrent networks have been used in [37] to accomplish this task. Dance-to-music involves creating a piece of music from a given dance [20, 27, 64]. For example, Zhun et al. [64] utilizes a generative adversarial network to generate music from dance videos. On the other hand, motion generation involves generating diverse human motions using multimodal inputs, such as text [44, 56, 15, 5, 57], music [27, 18, 52, 42] and incomplete motion [31, 1, 3]. Text-to-motion is one of the most important motion generation tasks. Recent works typically map text to motion using different architectures: diffusion model [57] and autoregressive transformer model [15]. Music-to-dance focuses on generating dance movements from music. For example, [42] predicts discrete token sequences conditioned on music, which are then used to regenerate the dance sequence. Motion Completion generates motion conditioning on partial motions, such as motion prediction [31, 1] and motion-in-between [34, 43]. Although these methods have shown promising results in various human motion tasks, most are limited to handling a single task. Until recently, some works [12, 21, 60, 62] attempt to integrate two or more tasks into a unified model, as shown in Tab. 3. However, these works either lack the ability of motion comprehension [62, 60] or fail to handle music modality [12, 21]. In this work, we propose a unified motion comprehension and generation framework that can handle multiple control signals simultaneously.
Language Models and Multimodal.  Large language models (LLMs) enabled by extensive datasets and model size, such as T5 [39], Flan-T5 [7], LLaMA [46], LLaMA-2 [47] and Vicuna [6], have demonstrated impressive comprehension and generation capabilities. Researchers have leveraged the capabilities of LLMs to handle multimodal tasks, expanding them to multimodal large language models (MLLMs). For example, AnyGPT [53] employs LLaMA-2 [47] to construct an any-to-any multimodal language model. NExT-GPT [50] employs Vicuna [6] with multimodal adaptors and diffusion decoders to perform tasks across arbitrary combinations of text, images, videos, and audio. Recently, the works [21, 60] attempt to use LLMs for motion-related tasks. [60] uses LLaMA [46] to build a general-purpose motion generator, which, however, lacks the ability to comprehend motion. [21] leverages T5 to construct a unified motion-language model, but cannot deal with music modality.
3 Method
To enhance the comprehension and generation of motion-relevant modalities, we propose a unified multimodal framework named M3GPT. As illustrated in Fig. 2, M3GPT consists of multimodal tokenizers responsible for compressing raw motion and music data into discrete tokens (Sec. 3.1), and a motion-aware language model that learns to understand and generate motion tokens from LLMs by corresponding text and music (Sec. 3.2). To address motion-relevant tasks, we employ a three-stage training scheme encompassing multimodal tokenizers training, modality-alignment pre-training, and instruction tuning (Sec. 3.3). During the inference process, multimodal tokens are decoded back into their original representations by associated de-tokenizers (decoders of multimodal tokenizers), enabling various motion-relevant tasks to be executed via instructions (Sec. 3.4).
Figure 2: An overview of the M3GPT framework. M3GPT consists of multimodal tokenizers and a motion-aware language model. The training process of M3GPT consists of three stages: multimodal tokenizers training, modality-alignment pre-training, and instruction tuning.
3.1 Multimodal tokenizers
As shown in Fig. 2, Multimodal tokenizers aim to discretize continuous human motion and music into language-like tokens, allowing the three modalities to be unified within a single language model.
3D Human Motion Tokenizer.  To represent motion in discrete semantic tokens, we build a 3D human motion tokenizer based on Vector Quantized Variational Autoencoders (VQ-VAE) following [12, 62, 21, 60]. The motion tokenizer consists of a motion encoder $\mathcal{E}_{m}$ and a motion decoder $\mathcal{D}_{m}$, along with a codebook $\mathcal{B}_{m}=\left\{b^{1},b^{2},\dots,b^{N_{m}}\right\}$ containing $N_{m}$ discrete semantic vectors. Notably, to facilitate mutual enhancement between motion and dance data, we employ a shared tokenizer for both motions and dances, projecting them into a consistent and shared semantic space. Formally, given a 3D motion sequence $\boldsymbol{m}\in\mathbb{R}^{T_{m}\times d_{m}}$, where $T_{m}$ is the time length and $d_{m}$ is the dimensionality of each frame’s pose, the motion encoder $\mathcal{E}_{m}$ that consists of several 1-D convolutional layers projects $\boldsymbol{m}$ to a latent embeddings $\boldsymbol{z}\in\mathbb{R}^{L_{m}\times d}$. Here, $L_{m}$ is the time interval after downsampling and $d$ is the latent dimension. Next, we transform $\boldsymbol{z}$ into a collection of codebook entries through discrete quantization. Specifically, the quantization process replaces each item of $\boldsymbol{z}$ with its nearest embedding in the codebook $\mathcal{B}_{m}$, obtaining the quantized latent vectors $\boldsymbol{e}\in\mathbb{R}^{L_{m}\times d}$ as follows:
$\boldsymbol{e}=\mathop{\arg\min}\limits_{b^{k}\in\mathcal{B}_{m}}\left\|\boldsymbol{z}-b^{k}\right\|_{2}.$ (1)
The motion decoder $\mathcal{D}_{m}$, which consists of several 1-D deconvolutional layers, projects the quantized embeddings back to raw motion space as $\boldsymbol{\hat{m}}=\mathcal{D}_{m}\left(\boldsymbol{e}\right)$. Following [21, 60], the motion tokenizer can be trained by the reconstruction loss, embedding loss and commitment loss as follows:
$\mathcal{L}_{vq}=\|\boldsymbol{\hat{m}}-\boldsymbol{m}\|_{1}+\|\mathrm{sg}\left[\boldsymbol{z}\right]-\boldsymbol{e}\|_{2}^{2}+\beta\|\boldsymbol{z}-\mathrm{sg}\left[\boldsymbol{e}\right]\|_{2}^{2}~{}.$ (2)
where $\mathrm{sg}\left[\cdot\right]$ is the stop gradient, and $\beta$ is the factor that adjusts the weight of the commitment loss.
After training the motion tokenizer, a motion sequence $\boldsymbol{m}$ can be represented as a sequence of discrete codebook-indices of quantized embedding vector, namely motion tokens $\boldsymbol{q_{m}}\in\mathbb{R}^{L_{m}}$, as follows:
$\boldsymbol{q_{m}}=\mathop{\arg\min}\limits_{k\in\left\{1,\dots,N_{m}\right\}}\left\|\mathcal{E}_{m}\left(\boldsymbol{m}\right)-b^{k}\right\|_{2}.$ (3)
Music Tokenizer.  For the music data, we adopt the VQ-VAE in Jukebox [9] as the music tokenizer, which consists of a music encoder $\mathcal{E}_{a}$, a music decoder $\mathcal{D}_{a}$ and a music codebook $\mathcal{B}_{a}$. Notably, the limited number of music samples in dance datasets makes it inadequate for training an effective music tokenizer. To leverage the strong representation ability of the tokenizer trained on the large-scale musical dataset, we use the pre-trained VQ-VAE from Jukebox [9], which has been trained on a dataset of 1.2 million songs. Specifically, we first segment each input music sample into 5-second music segments. Then, for each 5 seconds segment $\boldsymbol{a}\in\mathbb{R}^{T_{a}\times d_{a}}$, we use the pre-trained music tokenizer $\left\{\mathcal{E}_{a},\mathcal{B}_{a}\right\}$ to encode $\boldsymbol{a}$ into a sequence of discrete codebook-indices $\boldsymbol{q_{a}}\in\mathbb{R}^{L_{a}}$ (namely music tokens) following Eq. 3.
3.2 Language Model Backbone
Expanding Vocabulary.  To incorporate multimodal discrete representations into a pre-trained LLM, we expand the original text vocabulary $V_{t}$ in LLM with motion vocabulary $\mathcal{B}_{m}$ and music vocabulary $\mathcal{B}_{a}$, forming a new unified vocabulary $V=\left\{V_{t},\mathcal{B}_{m},\mathcal{B}_{a}\right\}$. To accommodate the expanded vocabulary, we extend the corresponding embedding and prediction layer of LLM, where the newly incorporated parameters are initialized randomly.
Unified Multimodal Language Model.  Equipped with multimodal tokenizers, we can compress multimodal data into discrete token sequences. To be specific, employing the trained motion tokenizer and music tokenizer, the input motion $\boldsymbol{m}\in\mathbb{R}^{T_{m}\times d_{m}}$ and music $\boldsymbol{a}\in\mathbb{R}^{T_{a}\times d_{a}}$ can be mapped into a sequence of discrete motion tokens $\boldsymbol{q_{m}}\in\mathbb{R}^{L_{m}}$ and music tokens $\boldsymbol{q_{a}}\in\mathbb{R}^{L_{a}}$. Then equipped with a unified vocabulary $V$, we can formulate various motion-relevant tasks in a general format, where both input and output tokens come from the same vocabulary. These tokens can represent natural language, human motion, music, or any combination, depending on the specific task to be solved. This naturally enables the core LLM to unify motion comprehension and generation tasks in an autoregressive manner.
Following [21], we employ T5 [39] as the language model backbone, which is pre-trained on 750 GB of text tokens. By leveraging this pre-trained large language model, we can harness its powerful modeling capabilities and generalizability to develop a more user-friendly, motion-related human-computer interaction model.
3.3 Training Strategy
The training process is divided into three stages. The first stage is Multimodal Tokenizers Training, which focuses on learning the motion/music tokenizer to represent motion/music as discrete tokens. The second stage is Modality-Alignment Pre-training, which aims to align motion, music, and text modalities, and facilitate collaboration across different motion-relevant tasks. The third stage is Instruction Fine-Tuning, aimed at enhancing the model’s instruction-following capability.
Stage1: Multimodal Tokenizers Training.  We first train a motion tokenizer using the objective defined in Eq. 2. As for the music tokenizer, due to the limited music samples in existing dance datasets, we directly use the pre-trained VQ-VAE model from Jukebox [9]. This process allows any motion sequence and music to be represented as a sequence of tokens, enabling seamless integration with text within LLM. To ensure the stability of LLM training, the encoder of motion tokenizer and whole music tokenizer remain unchanged. Notably, we continue to optimize the decoder of motion tokenizer in subsequent training stages to further enhance the quality of generated motions.
Stage2: Modality-Alignment Pre-training. To enable LLM to handle discrete modalities, we utilize paired motion corpus to train LLM in a next-token prediction task. This process aims to align the text, music, and motion modalities for unified reasoning in LLM.
  • Joint optimization of LLM and motion de-tokenizer.  Human motion (especially dance) encompasses intricate details. Previous works [21, 60] keep the motion de-tokenizer fixed during training LLM, which hinders LLM’s ability to perceive the distribution and details of motions. Specifically, in the output space of LLM, different motion tokens are treated as independent classes; therefore, the cost of classifying a motion token as semantic-similar token and semantic-distant token is the same. Apparently, relying solely on LLM’s autoregressive loss is insufficient for capturing the details of motion. To address this problem, we jointly optimize LLM and motion de-tokenizer in stage2 and stage3. This strategy enables the reconstruction error signals in raw motion space to backpropagate to LLM, enhancing LLM’s ability to generate the details of motion. With the goal of minimizing L1 loss between the predicted and real motion, we search for the motion’s token sequence that could minimize this L1 loss in original motion space. As the motion de-tokenizer continuously optimizes, the target motion’s token sequence, which supervises LLM training, dynamically changes. This dynamic adjustment reduces L1 loss progressively, achieving joint optimization.
  • Synergy learning of multitasks.  Although aligning text with one additional modality is relatively straightforward, integrating multiple modalities (e.g., motion, text, and music) within a single framework poses a significant challenge. Additionally, as noted in [4], multitask joint training usually achieves inferior performance on each individual task compared to single-task training. This phenomenon is also observed in our text-to-motion task, as shown in Tab. 2. We argue that the large modality difference among different motion-relevant tasks (e.g., music-to-dance and text-to-motion) prevents the model from effectively establishing connections between these tasks. Thus it is difficult for the model to identify a common optimization direction that benefits all tasks.
    As ‘text’ serves as a highly semantic descriptor for other modalities and is the most familiar and well-modeled modality to LLM, we use ‘text’ as a bridge to align motion, text, and music data, thereby mitigating conflicts in aligning multiple modalities. Initially, we construct paired textual descriptions for music samples in the dance datasets. Specifically, we use the style annotations of the music to create paired texts, such as ‘a person is dancing Jazz’. Then, we construct two auxiliary tasks using the generated pairs of music and text, i.e., music-to-text and text-to-dance. Through these two auxiliary tasks, M3GPT implicitly learns to decompose the complex music-to-dance task into two simpler tasks music-to-text and text-to-dance. Additionally, with a shared tokenizer for motion and dance, text-to-dance and text-to-motion tasks occupy the same matching space, and thus can mutually reinforce each other. In this way, M3GPT builds the synergies between the two primary motion generation tasks, music-to-dance and text-to-motion, facilitating mutual reinforcement, as shown in Tab. 2.
Combining the above analysis, we jointly train LLM and motion de-tokenizer using a mixture of motion comprehension and generation tasks, along with two auxiliary music-to-text and text-to-dance tasks. Besides the auxiliary tasks, we consider $2$ basic motion comprehension tasks, i.e., motion-to-text and dance-to-music, and $4$ basic motion generation tasks, i.e., text-to-motion, music-to-dance, motion prediction and motion in-between. Formally, for a specific task, we denote the source input consisting of a sequence of tokens as $\boldsymbol{q}_{s}=\left\{\boldsymbol{q}_{s}^{i}\right\}_{i=1}^{L_{s}}$, the target output as $\boldsymbol{q}_{t}=\left\{\boldsymbol{q}_{t}^{i}\right\}_{i=1}^{L_{t}}$, LLM predicts the probability distribution of potential next token at each step $p_{\theta}\left(\boldsymbol{q}_{t}^{i}|\boldsymbol{q}_{t}^{<i},\boldsymbol{q}_{s}\right)$ in an autoregressive manner. For motion generation tasks, we add a reconstruction loss. Specifically, when the output tokens are motion tokens, we pass them to motion de-tokenizer to generate a motion sequence (denoted as $\boldsymbol{\hat{m}}$), where a reconstruction loss is then employed for guidance. Overall, during this training process, the objective is to maximize the log-likelihood of the data distribution and minimize the reconstruction error within raw motion space:
$\mathcal{L}=\sum_{i=0}^{L_{t}-1}\log p_{\theta}\left(\boldsymbol{q}_{t}^{i}|\boldsymbol{q}_{t}^{<i},\boldsymbol{q}_{s}\right)+\lambda\left\|\boldsymbol{\hat{m}}-\boldsymbol{m}\right\|_{1},$ (4)
where $\boldsymbol{m}$ denotes the ground-truth for $\hat{\boldsymbol{m}}$ generated by motion de-tokenizer, and $\lambda$ is a hyper-parameter to adjust the weight for reconstruction loss.
Stage3: Instruction Fine-Tuning.  To enhance the generalization and instruction-following capability of M3GPT, we construct a multimodal instruction dataset with resort to GPT-4, building upon existing motion datasets. Specifically, we define $11$ core tasks, each comprising $200/50/50$ training/validation/test instruction templates. For example, an instruction for text-to-motion task could be "Create a motion that complements the poetic elements in <Caption_Placeholder>", with <Caption_Placeholder> standing for any text sequence; an instruction for music-to-dance could be "Script a dance that adapts to the tempo shifts in <Audio_Placeholder>", with <Audio_Placeholder> standing for any music sequence. Further details are available in Appendix B.4.
3.4 Inference M3GPT
During inference, we evaluate M3GPT’s performance across multiple motion-relevant tasks and datasets (Sec. 4.2 and Appendix (C, D). Also, we consider two challenging dance generation tasks to evaluate the zero-shot generalization ability of M3GPT:
(1) Generating long-duration dances from long music.  Long-duration dance generation involves creating uninterrupted, coherent dance sequences based on a single piece of music. Due to the limitation of computational cost and memory overload, we train M3GPT on the task of 5-second music-to-dance generation. Conversely, during inference, we can combine this training task music-to-dance to generate an initial 5-second dance segment, and an unseen zero-shot task music+dance-to-dance that recursively generates subsequent dance segments conditioned on both music and previously generated dance segments, to perform long-duration and coherent dance generation.
(2) Generating dance controlled by both music and text.  Integrating music and text as control signals in dance generation (music+text-to-dance) augments music-to-dance task with text modality. This process guides the generated dances to synchronize with particular actions described in input texts. Thanks to the tokenizer mechanism, M3GPT can seamlessly combine music and text in LLM’s input, enabling the integration of text instructions to produce a wide variety of dance sequences.
4 Experiments
4.1 Experimental setup
Datasets and Preprocessing.  We use a large-scale text-to-motion dataset: Motion-X [29], and two music-to-dance datasets: AIST++ [24] and FineDance [25]. Notably, the 3D pose annotations differ among these datasets, therefore, we standardize their processing for uniform usage. Specifically, we select $22$ joints common to these datasets and preprocess the motion samples following [13], resulting in motion sequences with identical representation. Further details on datasets and preprocessing are provided in Appendix B.1.
Evaluation Metrics.  Different tasks employ distinct evaluation metrics. We use the most common evaluation metrics to assess the performance of M3GPT for each task. (1) Text-to-Motion. Following [21, 12], we use Frechet Inception Distance (FID), Diversity (Div), R-Precition that calculates the top-1 motion-to-text retrieval accuracy (R TOP1). (2) Motion-to-Text. Following [21], we use linguistic metrics like BLEU, CIDEr, along with R-Precision for evaluating motion-to-text alignment. (3) Music-to-Dance. Following [26, 52], we use FID, Diversity and Beat Align Score (BAS) on kinetic features [22] (denoted as "k") to evaluate the dance generation quality. Notably, as noted in [48], the geometric features [33] are unreliable as a measure of dance generation quality. So we only use the kinetic features for evaluation. (4) Dance-to-Music. Following [52], we use Beats Coverage Scores (BCS), Beats Hit Scores (BHS), and F1 score to evaluate music generation quality. (5) Motion Prediction and In-Between. Following [21], we use FID and Diversity to measure the consistency between the provided pose conditions and generated motion. More details and results on other evaluation metrics are provided in Appendix B.2 and C.
Implementation Details.  For motion tokenizer, we set the codebook size to $512$. As for music tokenizer, we use the pre-trained VQ-VAE from Jukebox with a codebook size of $2048$. In term of temporal downsampling rate, the motion encoder uses a rate of $4$, while the music encoder uses a rate of $128$. We utilize T5 base [39] as our language model backbone. For training the motion tokenizer, we use Adam as the optimizer with a batch size of $1000$ and an initial learning rate of $10^{-4}$. To train the language model backbone, we employ the Adafactor_dev optimizer and use CosineAnnealingLR as the learning rate scheduler. The learning rate is set to $2\times 10^{-4}$ for pre-training stage, and $10^{-4}$ for instruction fine-tuning stage. For hyperparameter settings, $\lambda$ in Eq. 4 is set to $0.2$, and $\beta$ in Eq. 2 is set to $0.02$. All our experiments are conducted on 8 NVIDIA A40 GPUs. To evaluate the model’s performance across different platforms, we also test our trained M3GPT with T5-base on Ascend 910B NPUs. Further details on implementation and hyperparameter analysis are provided in Appendix E.
Table 2: Evaluation of synergy learning and joint optimization of LLM and motion de-tokenizer on Text-to-Motion (Motion-X [29]) and Music-to-Dance (AIST++ [24]). T2M: Text-to-Motion. A2D: Music-to-Dance. T2D: Text-to-Dance. A2T: Music-to-Text. Trained single task refers to a model trained and tested on a single task. Pre-trained and Instruction-tuned indicate the model after pre-training (stage2) and instruction tuning (stage3), followed by direct testing on each task. The arrows ($\uparrow$) indicate that higher values are better. The arrows ($\downarrow$) indicate that smaller values are better. Bold indicates the best result.
Methods Re-Optimizing motion de-tokenizer Text-to-Motion Music-to-Dance
[…]
Ground Truth - 0.675 0.009 2.316 17.10 8.19 0.2374
Trained single task 0.645 0.081 2.124 83.33 5.18 0.1835
Trained single task 0.656 0.078 2.133 75.47 5.57 0.1884
T2M+A2D 0.564 0.094 2.080 51.26 6.73 0.2037
T2M+A2D 0.578 0.092 2.106 47.71 7.47 0.1958
T2M+A2D+T2D+A2T 0.617 0.093 2.110 42.70 7.54 0.2084
T2M+A2D+T2D+A2T 0.626 0.088 2.197 25.24 7.63 0.2217
M3GPT (Pretrained without T2D and A2T) 0.526 0.105 2.058 40.71 7.47 0.2030
M3GPT (Pretrained without T2D and A2T) 0.547 0.104 2.099 37.14 7.61 0.2005
M3GPT (Pre-trained) 0.598 0.089 2.218 32.71 7.43 0.2090
M3GPT (Pre-trained) 0.601 0.092 2.251 27.65 7.52 0.2105
M3GPT (Instruction-tuned) 0.606 0.091 2.251 28.46 7.49 0.2052
M3GPT (Instruction-tuned) 0.615 0.093 2.253 24.34 7.50 0.2056
Table 3: Comparison results on Motion-X [29] dataset. The evaluation metrics are computed using the encoder introduced in Appendix A. Empty columns of previous methods indicate that they can not handle the task. Instruction-tuned only T2M indicates the model that is initially pre-trained on multiple tasks, followed by instruction tuning solely on text-to-motion task.
Methods Text-to-Motion Motion-to-Text Motion Prediction Motion In-between
[…]
Real $\text{0.675}^{\pm\text{0.003}}$ $\text{0.009}^{\pm\text{0.000}}$ $\text{2.316}^{\pm\text{0.011}}$ 0.881 - - 0.009 2.316 0.009 2.316
MLD [5] $\text{0.612}^{\pm\text{0.003}}$ $\text{0.122}^{\pm\text{0.008}}$ $\text{2.267}^{\pm\text{0.018}}$
T2M-GPT [55] $\text{0.647}^{\pm\text{0.002}}$ $\text{0.101}^{\pm\text{0.005}}$ $\underline{\text{2.270}}^{\pm\text{0.033}}$ 0.814 1.755 - -
MotionDiffuse [57] $\text{0.659}^{\pm\text{0.002}}$ $\underline{\text{0.075}}^{\pm\text{0.004}}$ $\text{2.220}^{\pm\text{0.022}}$
TM2T [15] $\text{0.581}^{\pm\text{0.002}}$ $\text{0.148}^{\pm\text{0.003}}$ $\text{2.005}^{\pm\text{0.034}}$ 0.806 12.13 20.16 - - - -
MDM [44] $\text{0.472}^{\pm\text{0.008}}$ $\text{0.078}^{\pm\text{0.000}}$ $\text{2.133}^{\pm\text{0.012}}$ 1.028 1.746 0.831 1.768
MoMask [16] $\textbf{\text{0.668}}^{\pm\text{0.003}}$ $\textbf{\text{0.074}}^{\pm\text{0.004}}$ $\text{2.241}^{\pm\text{0.016}}$ 0.626 1.884
MotionLCM [8] $\text{0.658}^{\pm\text{0.005}}$ $\text{0.078}^{\pm\text{0.003}}$ $\text{2.206}^{\pm\text{0.026}}$
MotionGPT[21] $\text{0.659}^{\pm\text{0.003}}$ $\text{0.078}^{\pm\text{0.001}}$ $\text{2.166}^{\pm\text{0.026}}$ $\underline{\text{0.840}}$ 11.21 31.36 0.701 1.818 0.648 1.875
Trained single task $\text{\text{0.656}}^{\pm\text{0.002}}$ $\text{0.078}^{\pm\text{0.002}}$ $\text{2.133}^{\pm\text{0.012}}$ 0.767 10.14 22.92 0.774 1.778 0.692 1.810
M3GPT (Pre-trained) $\text{0.601}^{\pm\text{0.002}}$ $\text{0.092}^{\pm\text{0.002}}$ $\text{2.251}^{\pm\text{0.012}}$ 0.834 11.00 24.12 0.707 1.874 0.604 1.879
M3GPT (Instruction-tuned) $\text{\text{0.615}}^{\pm\text{0.003}}$ $\text{0.093}^{\pm\text{0.002}}$ $\text{2.253}^{\pm\text{0.026}}$ 0.845 11.50 42.93 0.682 1.838 0.612 1.900
M3GPT (Instruction-tuned only T2M) $\underline{\text{0.661}}^{\pm\text{0.003}}$ $\text{0.076}^{\pm\text{0.002}}$ $\textbf{\text{2.273}}^{\pm\text{0.026}}$
Table 4: Comparison results on Motion-X [29] dataset based on Ascend 910B NPUs.
Methods Text-to-Motion Motion-to-Text Motion Prediction Motion In-between
[…]
Trained single task $\underline{\text{0.654}}^{\pm\text{0.002}}$ $\underline{\text{0.081}}^{\pm\text{0.003}}$ $\underline{2.304}^{\pm\text{0.017}}$ 0.763 10.16 22.89 0.776 1.818 0.712 1.880
M3GPT (Pre-trained) $\text{0.596}^{\pm\text{0.002}}$ $\text{0.096}^{\pm\text{0.003}}$ $\text{2.241}^{\pm\text{0.018}}$ 0.831 11.05 24.03 0.710 1.882 0.608 1.874
M3GPT (Instruction-tuned) $\text{\text{0.612}}^{\pm\text{0.002}}$ $\text{0.094}^{\pm\text{0.002}}$ $\text{2.276}^{\pm\text{0.021}}$ 0.846 11.52 42.64 0.684 1.841 0.621 1.903
M3GPT (Instruction-tuned only T2M) $\textbf{\text{0.659}}^{\pm\text{0.003}}$ $\textbf{\text{0.078}}^{\pm\text{0.003}}$ $\textbf{\text{2.314}}^{\pm\text{0.023}}$
Table 5: Comparison results on AIST++ [24] and FineDance [25] datasets.
Methods Music-to-Dance on AIST++ Music-to-Dance on FineDance Dance-to-Music on AIST++
[…]
Real 17.10 10.60 0.2374 - - 0.2120 - - -
FACT [24] 35.35 5.94 0.2209 113.38 3.36 0.1831 - - -
Bailando [42] 28.16 $\underline{\text{7.83}}$ 0.2332 82.81 7.74 0.2029 - - -
EDGE [48] 42.16 3.96 0.2334 94.34 8.13 0.2116 - - -
Lodge [26] 37.09 5.58 0.2423 45.56 6.75 0.2397 - - -
Foley [11] - - - - - - 96.4 41.0 57.5
CMT [10] - - - - - - 97.1 46.2 62.6
D2MGAN [64] - - - - - - 95.6 88.7 93.1
CDCD [65] - - - - - - 96.5 89.3 92.7
LORIS [52] - - - - - - 98.6 90.8 94.5
Trained single task 75.47 5.57 0.1884 128.37 6.48 0.2036 93.9 93.6 92.8
M3GPT (Pre-trained) 27.65 7.52 0.2105 92.35 7.67 0.2134 93.4 93.8 94.2
M3GPT (Instruction-tuned) $\underline{\text{24.34}}$ 7.50 0.2056 86.47 7.75 0.2158 93.6 94.0 94.9
M3GPT (Instruction-tuned for single task) 23.01 7.85 0.2261 42.66 8.24 0.2231 94.3 94.0 95.0
Table 6: Comparison results on AIST++ [24] and FineDance [25] datasets based on Ascend 910B NPUs.
Methods Music-to-Dance on AIST++ Music-to-Dance on FineDance Dance-to-Music on AIST++
[…]
Trained single task 77.32 5.61 0.1860 134.66 6.52 0.2088 93.8 93.6 92.7
M3GPT (Pre-trained) 27.99 $\underline{\text{7.61}}$ $\underline{\text{0.2102}}$ 91.39 7.71 0.2128 93.4 93.6 94.2
M3GPT (Instruction-tuned) $\underline{\text{25.05}}$ 7.48 0.2072 88.25 7.76 0.2160 93.5 93.8 94.7
M3GPT (Instruction-tuned for single task) 23.68 7.83 0.2264 43.78 8.39 0.2225 94.3 94.0 94.9
4.2 Ablation Studies
In this section, we conduct ablation studies to validate the effectiveness of our method. We use the same model architecture throughout the experiments. The ablation results are shown in Tab. 2.
Effectiveness of joint optimization of LLM and motion de-tokenizer.  Different from previous works [21, 60] that fix motion de-tokenizer during training LLM, we jointly optimize LLM and motion de-tokenizer in stage2 and stage3, as detailed in Sec. 3.3. As shown in Tab. 2, the joint optimization consistently brings performance gains across various metrics and most settings. Specifically, it largely enhances the fidelity of generated dances, reflected in a notable decrease in FIDk score. We also notice a minor increase (less than $0.003$) in FID of text-to-motion task in M3GPT. The possible reason is that the motion patterns controlled by text are relatively simple, making LLM optimized solely in discrete semantic space adequate for text-to-motion. Conversely, dances involve greater complexity, necessitating the joint optimization of motion decoder to accurately capture intricate dance movements without compromising information.
Effectiveness of synergy learning of multitasks.  During the training of M3GPT, we introduce a synergy multitask learning strategy by constructing two auxiliary tasks: Text-to-Dance (T2D) and Music-to-Text (A2T), as detailed in Sec. 3.3. As shown in Tab. 2, the inclusion of T2D and A2T consistently brings performance gains across various metrics on both text-to-motion and music-to-dance tasks. Specifically, for music-to-dance, the FIDk score is decreased by nearly $10$ points, indicating that the synergy learning helps generate more realistic dances. We argue that by incorporating these two auxilary tasks, M3GPT implicitly learns to decompose the complex music-to-dance into two simpler tasks, music-to-text and text-to-dance. This way, the text-to-motion task can assist in learning the music-to-dance task, thereby enhancing its performance.
4.3 Comparisons with State-of-the-arts
In this section, we compare our M3GPT with state-of-the-arts on multiple core motion-relevant tasks. We respectively report the comparison results on text-to-motion dataset, Motion-X [29], and music-to-dance datasets, AIST++ [24] and FineDance [25]. More quantitative and qualitative results are provided in Appendix C and D. Also, in the supplementary material’s zip file, we provide the render videos of generated motions/dances and generated music files by our M3GPT.
Main results on text-to-motion dataset.  On the text-to-motion dataset, Motion-X, we evaluate M3GPT on $4$ tasks, i.e., text-to-motion, motion-to-text, motion prediction, and motion in-between. The comparison results are shown in Tab. 6. As shown, M3GPT achieves competitive performance across all evaluated tasks, highlighting its capability to address diverse motion tasks in a single model. Also, for text-to-motion task, M3GPT (instruction-tuned only T2M), which combines multitask pre-training and instruction fine-tuning solely on T2M task, yields better performance than Trained single task that only trains the model on T2M task. Tab. 6 presents the results of testing on the Ascend 910B NPUs, where M3GPT also achieves comparably good performance. These results demonstrate that multitask pre-training can enhance the performance of individual tasks across different computation platforms. Further results of M3GPT trained on NPUs will be presented later.
Main results on music-to-dance datasets.  On the music-to-dance datasets, AIST++ and FineDance, we evaluate M3GPT on $2$ tasks, i.e., music-to-dance and dance-to-music. As shown in Tab. 6, in general, the performance of multitask pre-training and instruction fine-tuning in M3GPT outperforms single-task training, underscoring the effectiveness of multitask training for dance-relevant tasks. Also, M3GPT achieves competitive performance on most metrics. For music-to-dance, the best-performing method on the FineDance dataset is Lodge [26]. This approach features a specialized two-stage architecture for generating long-duration dance sequences, progressively refining movements from coarse to fine granularity using a diffusion model. On AIST++ dataset, M3GPT reports the best FIDk of $24.34$ for music-to-dance task, the best BHS and F1 of $94.0$ and $94.9$ for dance-to-music task. The results in Tab. 6, tested on the Ascend 910B NPUs, also demonstrate that multitask training can enhance the performance of both music-to-dance and dance-to-music tasks.
Figure 3: Qualitative results for long-term dance and music-text conditioned dance generation of M3GPT.
4.4 Evaluation on Zero-Shot Tasks
In this section, we explore M3GPT’s capabilities in handling zero-shot tasks, as mentioned in Sec. 3.4. Fig. 3 (a) shows the long-term dance generation. As seen, M3GPT can generate a coherent dance sequence by seamlessly integrating the music-to-dance and zero-shot music+dance-to-dance tasks. Fig. 3 (b) shows the generated 3D dance with both music and text instruction. We can see that M3GPT maintains plausible visual results in accordance with input text instructions (cartweel), underscoring its remarkable zero-shot generalization capability.
5 Conclusion
In this paper, we present M3GPT, a unified framework for comprehending and generating motion aligned with both text and music modalities. By employing text as a bridge, we build connections and synergies between different motion-relevant tasks, facilitating mutual reinforcement. Additionally, the joint optimization of LLM and motion de-tokenizer further enriches the details of generated motion, enhancing overall motion generation quality. Our extensive evaluations across various motion-relevant tasks demonstrate the effectiveness of M3GPT in both motion comprehension and generation. Besides, M3GPT exhibits strong zero-shot generalization abilities, enabling it to handle previously unseen and challenging motion-relevant tasks.
Limitations and Broader Impacts.  Although our M3GPT has successfully processed signals from motion, text, and music modalities for motion comprehension and generation, it focuses on modeling human body movements, excluding hands and faces details. Future research can extend the scope of M3GPT to include hands and facial modeling.
Acknowledgments. This work is sponsored by the National Natural Science Foundation of China (NSFC) under Grant 62306301, the National Postdoctoral Program for Innovative Talents under Grant BX20220310, and the CAAI-CANN Open Fund developed on the OpenI Community. It is also supported by the project of Peng Cheng Laboratory (PCL2023A08).
Appendix
Appendix A Text-Motion Alignment Model
Due to the lack of a powerful and publicly available text-motion alignment model, we independently leverage existing datasets to develop a functional text-motion alignment model, which is used to evaluate tasks such as text-to-motion, motion-to-text, motion prediction, and motion in-between. We adopt the motion encoder architecture $E_{m}$ from [14] and use the pretrained CLIP [38] ViT-B/32 model for the text encoder $E_{t}$. As depicted in Fig. 4, the training of the text-motion alignment model is split into two phases: pre-training the motion auto-encoder and text-motion contrastive learning. During the pre-training phase, we use motion data from the text-to-motion, and dance data from the music-to-dance tasks. This stage employs a reconstruction loss to ensure the model achieves a robust initial state capable of extracting an expressive motion representation. In the text-motion contrastive learning phase, we utilize text-motion pair data from the text-to-motion task. We incorporate an adapter MLP layer into both the motion and text encoders to align the dimensions of $\mathbf{z}_{m}$ and $\mathbf{z}_{t}$ at 512. This setup facilitates the alignment of text and motion in the representational space. The motion reconstruction loss $L_{\mathrm{recon\_motion}}$ for the pre-training stage and the contrastive learning loss $L_{\mathrm{infoNCE}}$ for the aligning stage are used to optimize this model, as follows,
$\mathcal{L}_{\mathrm{recon\_motion}}=\|x-\tilde{x}\|^{2}$ (5)
$\mathcal{L}_{\mathrm{infoNCE}}=-\frac{1}{N}\sum_{i=1}^{K}\log\left(\frac{\exp(\langle z_{i}^{\prime},z_{i}^{t}\rangle/\tau)}{\sum_{j=1}^{K}\exp(\langle z_{i}^{\prime},z_{j}^{t}\rangle/\tau)}\right)$ (6)
Figure 4: Pipeline of Text-Motion Alignment Model. The training of the text-motion alignment model includes two stages: pre-training motion auto-encoder and text-motion contrastive learning.
Appendix B Details for Training and Evaluating
B.1 Data Introductions and Preprocessing
We leverage the largest available text-to-motion dataset, Motion-X [29], along with widely-used music-to-dance datasets, AIST++ [24] and FineDance [25], for our multitask training regimen. Motion-X is used for text-to-motion, motion-to-text, motion prediction, and motion in-between tasks, while AIST++ and FineDance datasets support both music-to-dance and dance-to-music tasks, and are also adapted for motion prediction and in-between tasks to enhance our training resources.
Motion-X includes 15.6 million precise 3D whole-body SMPL-X [36] pose annotations across 81.1K motion sequences with sequence-level semantic text descriptions. AIST++ contains 1,409 dance motion sequences across 10 dance genres with SMPL [30] pose annotations, and FineDance provides 7.7 hours of dance, totaling 831,600 frames with SMPL-H [41] poses at 30 fps across 16 dance genres. Tab. 7 shows the training datasets and their corresponding sample numbers that we use to train our model.
We standardize data across these datasets by selecting the 22 common joints and normalizing each motion sequence to face the same direction and to run at 30 fps. We use a processing technique consistent with prior research [14, 5, 21] that integrates joint velocities, positions, and rotations for consistent motion representation, facilitating effective utilization across tasks. For the music data, we use the Librosa toolkit [32] to load raw .wav data at a sampling rate of 22050 Hz, processed into features by the Jukebox encoder [9]. To optimize the use of these datasets, we strategically employ data from specific datasets for different tasks. During training for music and dance tasks, we randomly select 5-second segments from complete music tracks and corresponding dance segments as training samples, setting the sample length to 6.25 seconds for motion prediction and in-between tasks with AIST++ and FineDance. When assessing music-to-dance on the FineDance dataset, we don’t evaluate all 5-second samples directly. Instead, we generate continuous long dance sequences using music-to-dance and motion prediction, then segment these into 30-second samples for evaluation to align with Lodge’s testing methodology.
Table 7: The training datasets and sample numbers for different tasks.
Tasks T2M M2T Motion Prediction/In-between A2D D2A A2T T2D
Training dataset Motion-X Motion-X Motion-X/AIST++/FineDance AIST++/FineDance AIST++/FineDance AIST++/FineDance AIST++/FineDance
Training samples number 64867 64867 64867/952/177 952/177 952/177 952/177 952/177
B.2 Comprehensive Evaluation Metrics
Different tasks utilize specific evaluation metrics. We use the consistent evaluation metrics following prior research [14, 21, 12, 26, 64, 52].
  • Text-to-Motion. We measure the discrepancy between generated and actual motion features using Frechet Inception Distance (FID), assess diversity (Div) among all generated motion sequences, and evaluate motion-text semantic correlation with R-precision. Multimodal Distance (MM Dist) quantifies the disparity between motions and texts. A specialized model developed for evaluating the text-to-motion task on the Motion-X dataset with 22 joints is detailed in Appendix A.
  • Motion-to-Text. We use linguistic metrics including BLEU [35], ROUGE [28], CIDEr [49], and BertScore [59], along with R-Precision and MM Dist to assess alignment between generated text and motion.
  • Music-to-Dance. We employ the evaluation framework recommended by FACT [24] and Bailando [42], utilizing FID, Diversity, and Beat Align Score (BAS) to gauge dance quality. In our paper, we use kinetic features to compute FID and Diversity.
  • Dance-to-Music. We use metrics from [52, 64] such as Beats Coverage Scores (BCS), Beats Hit Scores (BHS), F1 scores, Beats Coverage Std (CSD), and Beats Hit Std (HSD) to evaluate music generation quality.
  • Motion Prediction and Motion In-between. We use Average Displacement Error (ADE) and Final Displacement Error (FDE) to assess the quality of predicted motion. The text-motion alignment model aids in evaluating motion prediction performance.
B.3 Distributed Training for Multitasks
We employ a single-node multi-GPU distributed training strategy for M3GPT, distributing each task across separate GPUs to facilitate multitask training through shared model parameters. This method allows us to tailor the maximum token length of the LLM for each task, based on the longest sample sequence typical for that task. Specifically, we set the maximum LLM token length to $192$ for tasks such as text-to-motion, motion-to-text, motion prediction, and motion in-between. For tasks involving the music modality, such as music-to-dance and dance-to-music, the maximum length is set at $980$. This task-specific configuration enables us to optimize batch sizes effectively, thus maximizing GPU utilization.
In our experimental setup, the batch size is set to $40$ for the text-to-motion, text-to-dance, and motion-to-text tasks, and $4$ for the music-to-dance, dance-to-music, and music-to-text tasks. For motion prediction and motion in-between tasks, the batch size is set to $45$. We also establish the number of iterations for pre-training at 1,000,000, instruction fine-tuning at 200,000. This structured approach ensures that each task is optimally processed, enhancing the efficiency and effectiveness of our training regimen.
B.4 Tasks for Pre-training and Instruction tuning
As shown in Fig. 5, we define 11 core motion-related tasks for the instruction tuning of M3GPT, including text-to-motion, random-to-motion, motion-to-text, random-to-text, motion prediction, motion in-between, music-to-dance, dance-to-music, random-to-music, text-to-dance, and music-to-text. The tasks of text-to-dance and music-to-text were specifically constructed based on the music-to-dance datasets. Random-to-X represents the unconstrained generation of motion, text and music. In Tab. 8, we present a selection of command templates for each task, where <Motion_Placeholder>, <Caption_Placeholder>, and <Music_Placeholder> respectively represent the motion sequence (including dance sequence), textual description, and music segment from the training data.
Figure 5: Tasks for M3GPT pre-training and instruction tuning. Random represents the unconstrained generation of motion/text/music in the corresponding task.
Table 8: Examples of instruction templates for each task when instruction tuning M3GPT.
Task Instruction template Output
text-to-motion Design a motion that illustrates the emotions conveyed by <Caption_Placeholder>. <Motion_Placeholder>
How could you express the resilience mentioned in <Caption_Placeholder> through motion?
Can you develop a motion that captures the existential debate in <Caption_Placeholder>?
random-to-motion Can you generate a motion randomly? <Motion_Placeholder>
Please generate a random motion.
Display a motion for me.
motion-to-text What themes are explored through the motion in <Motion_Placeholder>? <Caption_Placeholder>
Can you describe the motion <Motion_Placeholder> with texts?
How would you interpret the actions depicted in <Motion_Placeholder>?
random-to-text Can you generate a text description for motion randomly? <Caption_Placeholder>
Give me a caption which describes a action.
How can we describe a motion with texts?
motion prediction What movements would suitably follow the thematic climax of <Motion_Placeholder>? <Motion_Placeholder>
What steps might deepen the emotional expression seen in <Motion_Placeholder>?
What movements could follow to resolve the suspense built in <Motion_Placeholder>?
motion in-between What new character dynamics could the middle section of <Motion_Placeholder> explore? <Motion_Placeholder>
Infer the type of dramatic climax that the masked section of <Motion_Placeholder> might contain.
What potential themes of ascent or descent could be explored in the middle of <Motion_Placeholder>?
music-to-dance Script a dance that adapts to the tempo shifts in <Music_Placeholder>. <Motion_Placeholder>
Create a dance that would visually mimic the lyrical journey in <Music_Placeholder>.
Compose a dance that explores the genre characteristics of <Music_Placeholder>.
dance-to-music Can you design a music for this dance <Motion_Placeholder>? <Music_Placeholder>
Please create a music based on this dance <Motion_Placeholder>.
Arrange a symphony that captures the shifts in <Motion_Placeholder>.
random-to-music Please generate a music for a dance randomly. <Music_Placeholder>
Can you generate a music with dance style?
Creat a music for any style dance.
text-to-dance Please generate a dance based on the caption <Caption_Placeholder>. <Motion_Placeholder>
Create a dance for the text <Caption_Placeholder>.
Generate a dance that corresponds to the textual description <Caption_Placeholder>.
music-to-text Describe the dance movements that correspond to the given music <Music_Placeholder>. <Caption_Placeholder>
Describe the dance actions that match the provided music <Music_Placeholder>.
Detail the dance steps associated with the specified music <Music_Placeholder>.
Appendix C Quantitative Results and Comparisons with SOTA Methods
In this section, we compare the performance of our M3GPT with existing SOTA methods on a broader set of metrics for each task across three datasets: Motion-X [29], AIST++ [24], and FineDance [25]. Tab. 12 shows the comparison of text-to-motion on Motion-X dataset. Tab. 12 shows the comparison of motion-to-text on Motion-X dataset. Tab. 12 shows the comparison of music-to-dance on AIST++ and FineDance datasets. Tab. 12 shows the comparison of dance-to-music on AIST++ and FineDance datasets. Tab. 18 and Tab. 18 shows the comparison of motion prediction and motion in-between on Motion-X dataset. Tab. 18 shows the comparison of music-to-text, text-to-dance and text-to-music on AIST++ dataset. Tab. 18 and Tab. 18 show the comparison of motion-related tasks among different size of T5. As shown in these tables, M3GPT can achieve competitive performance with SOTAs across all evaluated tasks.
Appendix D Qualitative Results and Comparison with SOTA Methods
Fig. 6 presents visualizations for a variety of tasks, including text-to-motion, motion-to-text, motion prediction, motion in-between, music-to-dance, long-term dance generation, and music-text conditioned dance generation. The visualization results show that our method can generate realistic results across various motion-relevant tasks. Fig. 7 presents the qualitative results between different methods for text-to-motion and music-to-dance.
Appendix E Additional Experiments
The performance of text-to-motion based on different $\lambda$.  In Tab. 18, the performance of M3GPT on Motion-X dataset is analyzed across different values of $\lambda$ for the text-to-motion task. The results indicate that as $\lambda$ increases, the model’s recall precision (Top1, Top2, Top3) initially improves but subsequently declines. The optimal performance is achieved at $\lambda=0.2$, as evidenced by the highest scores in RPrecision and modality metrics. Further increases in $\lambda$ to 0.3 and 0.4 lead to deteriorating performance, particularly in FID and RPrecision, suggesting that excessive $\lambda$ values may result in over-regularization or reduced adaptability of the model.
Table 9: Comparison of Text-to-Motion on Motion-X dataset [29]. The arrows ($\uparrow$) indicate that higher values are better. The arrows ($\downarrow$) indicate that smaller values are better. Bold and underline indicate the best and the second best result.
Methods RPrecision $\uparrow$ FID $\downarrow$ MM Dist $\downarrow$ Div $\uparrow$ MModality $\uparrow$
Top1 Top2 Top3
Real $\text{0.675}^{\pm\text{0.003}}$ $\text{0.821}^{\pm\text{0.003}}$ $\text{0.878}^{\pm\text{0.002}}$ $\text{0.009}^{\pm\text{0.000}}$ $\text{2.938}^{\pm\text{0.007}}$ $\text{2.316}^{\pm\text{0.011}}$
MDM [44] $\text{0.472}^{\pm\text{0.008}}$ $\text{0.616}^{\pm\text{0.005}}$ $\text{0.704}^{\pm\text{0.003}}$ $\text{0.161}^{\pm\text{0.006}}$ $\text{5.404}^{\pm\text{0.031}}$ $\text{2.234}^{\pm\text{0.015}}$ $\underline{\text{2.241}}^{\pm\text{0.043}}$
MLD [5] $\text{0.612}^{\pm\text{0.003}}$ $\text{0.743}^{\pm\text{0.002}}$ $\text{0.808}^{\pm\text{0.004}}$ $\text{0.122}^{\pm\text{0.008}}$ $\text{3.117}^{\pm\text{0.035}}$ $\text{2.267}^{\pm\text{0.018}}$ $\text{2.210}^{\pm\text{0.055}}$
T2M-GPT [55] $\text{0.647}^{\pm\text{0.002}}$ $\text{0.785}^{\pm\text{0.004}}$ $\text{0.845}^{\pm\text{0.003}}$ $\text{0.101}^{\pm\text{0.005}}$ $\text{3.046}^{\pm\text{0.028}}$ $\underline{\text{2.270}}^{\pm\text{0.033}}$ $\textbf{2.226}^{\pm\text{0.036}}$
MotionDiffuse [55] $\underline{\text{0.659}}^{\pm\text{0.002}}$ $\underline{\text{0.802}}^{\pm\text{0.004}}$ $\textbf{0.865}^{\pm\text{0.002}}$ $\textbf{0.075}^{\pm\text{0.004}}$ $\text{2.944}^{\pm\text{0.004}}$ $\text{2.220}^{\pm\text{0.022}}$ $\text{2.102}^{\pm\text{0.036}}$
Trained single task $\text{0.656}^{\pm\text{0.002}}$ $\text{0.795}^{\pm\text{0.001}}$ $\text{0.843}^{\pm\text{0.001}}$ $\text{0.078}^{\pm\text{0.000}}$ $\underline{\text{2.942}}^{\pm\text{0.001}}$ $\text{2.133}^{\pm\text{0.012}}$ $\text{2.046}^{\pm\text{0.052}}$
M3GPT (Pre-trained) $\text{0.601}^{\pm\text{0.002}}$ $\text{0.751}^{\pm\text{0.003}}$ $\text{0.803}^{\pm\text{0.002}}$ $\text{0.092}^{\pm\text{0.002}}$ $\text{2.945}^{\pm\text{0.001}}$ $\text{2.251}^{\pm\text{0.012}}$ $\text{2.188}^{\pm\text{0.074}}$
M3GPT (Fine-tuned) $\text{0.615}^{\pm\text{0.003}}$ $\text{0.757}^{\pm\text{0.004}}$ $\text{0.815}^{\pm\text{0.003}}$ $\text{0.093}^{\pm\text{0.002}}$ $\text{2.944}^{\pm\text{0.002}}$ $\text{2.253}^{\pm\text{0.023}}$ $\text{2.204}^{\pm\text{0.058}}$
M3GPT(Fine-tuned only T2M) $\textbf{0.661}^{\pm\text{0.003}}$ $\textbf{0.804}^{\pm\text{0.004}}$ $\underline{\text{0.861}}^{\pm\text{0.003}}$ $\underline{\text{0.076}}^{\pm\text{0.002}}$ $\textbf{2.940}^{\pm\text{0.002}}$ $\textbf{2.273}^{\pm\text{0.026}}$ $\text{2.131}^{\pm\text{0.032}}$
Table 10: Comparison of Motion-to-Text on Motion-X [29].
Methods RPrecision $\uparrow$ MM Dist $\downarrow$ Bleu@1 $\uparrow$ Bleu@4 $\uparrow$ Rouge $\uparrow$ CIDEr $\uparrow$ BertScore $\uparrow$
Top1 Top2 Top3
Real 0.681 0.824 0.881 2.897 - - - - -
TM2T [15] 0.574 0.726 0.806 2.988 30.54 12.13 32.52 20.16 25.37
Trained single task 0.565 0.706 0.767 3.011 31.07 10.14 31.65 22.92 28.19
M3GPT (Pre-trained) 0.627 0.773 0.834 2.946 33.31 11.00 34.10 24.12 30.96
M3GPT (Fine-tuned) 0.631 0.783 0.845 2.950 34.27 11.50 34.55 42.93 31.49
Table 11: Comparison of Music-to-Dance on AIST++ [24] and FineDance [25].
Methods Music-to-Dance on AIST++ Music-to-Dance on FineDance
[…]
Real 17.10 8.19 0.2374 - 9.73 0.2120
FACT [24] 35.35 5.94 0.2209 113.38 3.36 0.1831
Bailando [42] 28.16 7.83 0.2332 82.81 7.74 0.2029
EDGE [48] 42.16 3.96 0.2334 94.34 8.13 0.2116
Lodge [26] 37.09 5.58 0.2423 45.56 6.75 0.2397
Trained single task 75.47 5.57 0.1884 128.37 6.48 0.2036
M3GPT (Pre-trained) $\underline{\text{27.65}}$ $\underline{\text{7.52}}$ 0.2105 92.35 7.67 0.2134
M3GPT (Fine-tuned) 24.34 7.50 0.2056 86.47 7.75 0.2158
Table 12: Comparison of Dance-to-Music on AIST++ [24] and FineDance [25].
Methods Dance-to-Music on AIST++ Dance-to-Music on FineDance
BCS $\uparrow$ CSD $\downarrow$ BHS $\uparrow$ HSD $\downarrow$ F1 $\uparrow$ BCS $\uparrow$ CSD $\downarrow$ BHS $\uparrow$ HSD $\downarrow$ F1 $\uparrow$
Foley [11] 96.4 6.9 41.0 15.0 57.5 - - - - -
CMT [10] 97.1 6.4 46.2 18.6 62.6 - - - - -
D2MGAN [64] 95.6 9.4 88.7 19.0 93.1 - - - - -
CDCD [65] 96.5 9.1 89.3 18.1 92.7 - - - - -
LORIS [52] 98.6 6.1 90.8 13.9 94.5 - - - - -
Trained single task 93.9 9.2 93.6 12.8 92.8 84.84 21.61 51.35 27.13 63.97
M3GPT (Pre-trained) 93.4 10.9 93.8 11.5 94.2 83.16 19.95 73.65 23.90 78.12
M3GPT (Fine-tuned) 93.6 10.1 94.0 10.6 94.9 84.10 18.36 74.66 23.45 78.23
Table 13: Comparison of Motion Prediction and Motion In-between on Motion-X [29].
Methods Motion Prediction Motion In-between
FID $\downarrow$ Diversity $\uparrow$ ADE $\downarrow$ FDE $\downarrow$ FID $\downarrow$ Diversity $\uparrow$ ADE $\downarrow$
Ground Truth 0.009 2.316 - - 0.009 2.316 -
MDM [44] 1.028 1.746 8.057 11.266 0.831 1.768 6.542
Trained single task 0.774 1.778 7.840 9.575 0.692 1.810 6.690
M3GPT-pretrain 0.707 1.874 6.954 8.684 0.604 1.897 5.692
M3GPT-finetune 0.682 1.838 6.898 8.091 0.612 1.900 5.649
Table 14: Comparison of MPJPE on Motion Prediction and Motion In-between on Motion-X [29].
Methods on Motion-X Motion Prediction (MPJPE $\downarrow$) Motion In-between (MPJPE $\downarrow$)
T2M-GPT [55] 80.2 63.7
MoMask [16] 67.9 55.2
MotionGPT [21] 71.3 59.9
M3GPT (Instruction-tuned) 54.2 51.0
Table 15: Comparison of Music-to-Text (A2T), Text-to-Dance (T2D) and Text-to-Music (T2A) on AIST++.
Methods on AIST++ Music-to-Text Text-to-Dance Text-to-Music
Bleu@4 $\uparrow$ CIDEr $\uparrow$ R-TOP1 $\uparrow$ FID $\downarrow$ BCS $\uparrow$ BHS $\uparrow$
M3GPT (Single task training for A2T) 9.24 24.55 - - - -
M3GPT (Single task training for T2D) - - 0.541 0.095 - -
MusicLDM [2] - - - - 74.5 73.8
Mubert - - - - 73.3 73.0
M3GPT (Instruction-tuned) 11.95 28.88 0.588 0.077 74.5 74.7
Table 16: Comparison of Text-to-Motion and Motion-to-Text with different size of T5.
Methods on Motion-X LLM Training time Text-to-Motion Motion-to-Text
R-TOP1 $\uparrow$ FID $\downarrow$ Div $\uparrow$ R-TOP3 $\uparrow$ Bleu4 $\uparrow$ CIDEr $\uparrow$
M3GPT T5-small (60M) 5 days 0.598 0.096 2.202 0.822 10.43 38.22
M3GPT T5-base (220M) 7 days 0.615 0.093 2.253 0.845 11.50 42.93
M3GPT T5-large (770M) 8 days 0.619 0.090 2.256 0.848 11.64 43.05
Table 17: Comparison of Music-to-Dance and Dance-to-Music with different size of T5.
Methods on AIST++ LLM Training time Music-to-Dance Dance-to-Music
FIDk $\downarrow$ DIVk $\uparrow$ BAS $\uparrow$ BCS $\uparrow$ BHS $\uparrow$
M3GPT T5-small (60M) 5 days 28.05 5.96 0.2091 89.1 91.2
M3GPT T5-base (220M) 7 days 24.34 7.50 0.2056 93.6 94.0
M3GPT T5-large (770M) 8 days 23.26 7.54 0.2061 93.8 94.1
Table 18: Hyper-parameter analysis of $\lambda$. Comparison of Text-to-Motion on Motion-X [29] with different values of $\lambda$. For this ablation study, M3GPT is trained solely on the text-to-motion task to examine the impact of $\lambda$. This study is conducted during the pre-training stage.
Methods RPrecision $\uparrow$ FID $\downarrow$ MM Dist $\downarrow$ Diversity $\rightarrow$ MModality $\uparrow$
Top1 Top2 Top3
Real $\text{0.675}^{\pm\text{0.003}}$ $\text{0.821}^{\pm\text{0.003}}$ $\text{0.878}^{\pm\text{0.002}}$ $\text{0.009}^{\pm\text{0.000}}$ $\text{2.938}^{\pm\text{0.007}}$ $\text{2.316}^{\pm\text{0.011}}$
$\lambda$=0.0 $\text{0.645}^{\pm\text{0.002}}$ $\text{0.778}^{\pm\text{0.003}}$ $\text{0.826}^{\pm\text{0.002}}$ $\text{0.081}^{\pm\text{0.002}}$ $\text{2.944}^{\pm\text{0.002}}$ $\text{2.124}^{\pm\text{0.011}}$ $\text{2.025}^{\pm\text{0.021}}$
$\lambda$=0.1 $\text{0.649}^{\pm\text{0.002}}$ $\text{0.787}^{\pm\text{0.002}}$ $\text{0.838}^{\pm\text{0.002}}$ $\textbf{0.078}^{\pm\text{0.002}}$ $\text{2.944}^{\pm\text{0.003}}$ $\text{2.128}^{\pm\text{0.022}}$ $\text{2.039}^{\pm\text{0.049}}$
$\lambda$=0.2 $\textbf{0.656}^{\pm\text{0.002}}$ $\textbf{0.795}^{\pm\text{0.001}}$ $\textbf{0.843}^{\pm\text{0.001}}$ $\textbf{0.078}^{\pm\text{0.000}}$ $\text{2.942}^{\pm\text{0.001}}$ $\textbf{2.133}^{\pm\text{0.012}}$ $\textbf{2.046}^{\pm\text{0.052}}$
$\lambda$=0.3 $\text{0.629}^{\pm\text{0.002}}$ $\text{0.716}^{\pm\text{0.002}}$ $\text{0.784}^{\pm\text{0.001}}$ $\text{0.095}^{\pm\text{0.002}}$ $\textbf{2.941}^{\pm\text{0.001}}$ $\text{2.113}^{\pm\text{0.037}}$ $\text{2.036}^{\pm\text{0.041}}$
$\lambda$=0.4 $\text{0.573}^{\pm\text{0.001}}$ $\text{0.725}^{\pm\text{0.002}}$ $\text{0.793}^{\pm\text{0.002}}$ $\text{0.102}^{\pm\text{0.002}}$ $\text{2.942}^{\pm\text{0.002}}$ $\text{2.090}^{\pm\text{0.007}}$ $\text{2.030}^{\pm\text{0.086}}$
Figure 6: The qualitative results for different motion comprehension and generation tasks.
Figure 7: Qualitative comparisons for text-to-motion task and music-to-dance task. (a) refers to the qualitative comparison between Real, MDM, MoMask and M3GPT on text-to-motion task. The red words and boxes highlight the misaligned motions. The results demonstrate that our M3GPT shows good text understanding for motion generation. (b) refers to the qualitative comparison between Bailando and M3GPT on music-to-dance task. The input is an 5-second-long piece of music in the Break style.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: A summary of the paper’s contributions is provided in the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Not Applicable
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Section 4 and Appendix for implementation details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: See Section 4 and Appendix for implementation details.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [N/A]
  33. Justification: Not Applicable.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: See Section 4 and Appendix for implementation details.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [Yes]
  48. Justification: See Section 5
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: Not Applicable.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: See References.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: Not Applicable.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: Not Applicable.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: Not Applicable.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.